The market for electricity in Utopia is characterised by the following equations

Qd = 100-5P ; Qs = 50+5P


(b) Suppose the government of Utopia set a maximum price of R2.50 per unit of electricity. Explain and show in the graph drawn in (a) the effect of the government intervention in the electricity market.


1
Expert's answer
2021-11-07T19:43:17-0500

(b) R2.5

Qd=100-5*2.5=87.5

Qs=50+5*2.5=62.5

Shortage=Qd-Qs= 87.5-62.5= 5.5